OpenStudy (anonymous):

3x + 9+4x = 30 Can some one show me how to write this out in long form?

OpenStudy (anonymous):

x=3

OpenStudy (saifoo.khan):

7x = 30 - 9 7x = 21 x = 21/7 x = 3

OpenStudy (mimi_x3):

7x -21=0 7x = 21 x = 3

OpenStudy (saifoo.khan):

U copied me! >.<

OpenStudy (saifoo.khan):

you*

OpenStudy (mimi_x3):

LOL, hell no why would i want to copy you Everyone would get the same answer And my answer looks way~ BETTER xD

OpenStudy (anonymous):

How did we get 7?

OpenStudy (saifoo.khan):

haha, from which angle does that looks good? u missed a step.

OpenStudy (anonymous):

Okay that wat I thought Thanks

OpenStudy (mimi_x3):

well , my ones easier to look at it, I don't need to waste times to do useless steps

OpenStudy (mimi_x3):

3x + 4x = 7x*

OpenStudy (saifoo.khan):